A 6.89 kg box sled is coasting across frictionless ice at a speed of 2.87 m/s when a 22.3 kg package is dropped into it from above. What is the new speed of the sled?
Answer #1

Given that
     m1 = 6.89 kg
     v1 = 2.87 m/s
     m2 = 22.3 + 6.89 = 29.19 kg
     v2 = ?

From conservation of momentum we have
          Pi = Pf
      m1v1 = m2v2
==> v2 = m1v1 / m2
             = 6.89 * 2.87 / 29.19
             = 0.6774 m/s